A gun fires a 50gm bullet at a speed of 1000m/s. The gunner can exert a force of 180N against the gun. Determined the maximum number of bullets he can fire per minute

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
 
 
here, *The gun will exert an impulsive force on the man when it will be fired. * As force will be of impulsive nature I=MV * 50*10^-3*1000=50 N-s * the man can hold with a force of 180 N for 1 sec. * The force must be of impulsive nature. * no. of bullets=180/50=3.6
 
*hence in 1 minute = 3.6*60 = 216 bullets
 
 
Regards
Arun (askIITians forum expert)